OpenStudy (anonymous):

Express 72 degrees in radian measure

OpenStudy (anonymous):

the conversion factor is pi/90

OpenStudy (anonymous):

2.513 radians

OpenStudy (radar):

72 is to 360 as x is to 2Pi

OpenStudy (anonymous):

72*pi/90

OpenStudy (anonymous):

is 2.513

OpenStudy (radar):

\[2 \pi radians = 360^{o}\]

OpenStudy (radar):

72 degrees x ----------- = --------- 360 degrees 2 Pi 72 (2) (Pi) = 360 x x=452.389/360= 1.256 radians

OpenStudy (anonymous):

wait, oh sorry, pi/180 doin too much calculus

OpenStudy (radar):

babygurl, have you got it?

OpenStudy (anonymous):

theres a formula that i learned something like pi=d/r

OpenStudy (radar):

Radians=(Pi)(degrees/180 degrees)

OpenStudy (radar):

www.themathpage.com/atrig/radian-measure.htm that URL has good explanation